There are two states of economy denoted as State A and B. Primary Security A is priced to be $0.35 and risk free interest rate is 12%

(A) What is the price of primary security B

(B) If BMI's stock is forecasted to be either $95 if state A occurs and $65 if state B occurs. What will be the price your willing to pay for BMI stock today?
Answer #1

Solution:

Price of primary security A is $0.35

(A) Price of primary security B is (1-0.35)= $0.65

(B) Forecasted price of BMI stock = $95*0.35 + $65*0.65= $75.50

Price of BMI stock today = 75.50/(1+0.12)= $67.41